OpenStudy (anonymous):

Checking my answer: Find the equation of the line that is perpendicular to the tangent line to the curve, y=(3x+1)/(4x-2)

OpenStudy (anonymous):

at point (1,2)

OpenStudy (anonymous):

I used the quotient rule to get -10/(8x^2 -16x+4)

OpenStudy (anonymous):

Plugged in 1 to get 5/2

OpenStudy (anonymous):

Since it is perpendicular the slope is the negative reciprocal, so -2/5

OpenStudy (anonymous):

my final answer is y-2=(-2/5)(x-1)

hartnn (hartnn):

derivative i got is -5/2 not +5/2

OpenStudy (anonymous):

ah yes that's right

hartnn (hartnn):

isn't the first term of the denominator \(\Large 16x^2\)

hartnn (hartnn):

and not 8x^2

OpenStudy (anonymous):

well I did (4x-2)(4x-2)

hartnn (hartnn):

yeah, first term = 4x*4x = 16x^2

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

of course

hartnn (hartnn):

so slope = -5/2 :)

hartnn (hartnn):

y-2=(2/5)(x-1)

OpenStudy (anonymous):

Cool, I'll try to be more careful next time ^_^

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^